OpenStudy (anonymous):

what is 26.5% pf 46?

OpenStudy (anonymous):

x % of y is written as : \[\frac{x}{100} \times y\]

OpenStudy (anonymous):

So, 26.5% of 46 will be written as: \[\frac{26.5}{100} \times 46\] Solve it..

OpenStudy (anonymous):

12.19

OpenStudy (anonymous):

Just look it up on google lol

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

my name explains it all.

OpenStudy (anonymous):

you must read it @Caniff http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Don't just put here answers, write with proper reasoning.. It is website for Learning.. If you are to use Google, then why are you here??

OpenStudy (anonymous):

and the flame war begins on my question board!

OpenStudy (mrnood):

@Caniff From the code of conduct: Give Help, Not Answers –I will encourage and guide those needing help, and not just give them an answer Don't devalue the question/answer process! •Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. •Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!

OpenStudy (anonymous):

whoops, sorry

OpenStudy (anonymous):

Didnt know

OpenStudy (anonymous):

caniff its ok... thank you for answering me :P

OpenStudy (anonymous):

But thanks, i wont do that again

OpenStudy (anonymous):

No promblem

OpenStudy (anonymous):

@Caniff why can't you google for the question you have posted?? :P

OpenStudy (anonymous):

I did but i still dont understand

OpenStudy (anonymous):

Okay just kidding, nothing personal.. :)

OpenStudy (anonymous):

Okie Dokie

OpenStudy (anonymous):

;P